OpenStudy (anonymous):

evaluate 5a+5b where a=-6 and b=-5?

OpenStudy (anonymous):

So the equation is (5 x a) + (5 x b) Now replace a and b with the given values (5 x -6) + (5 x -5) (-30) + (-25) = -30 - 25 = -55

OpenStudy (anonymous):

thank u!!!